# Circuits and Backdoors: Five Shades of the SETH

### v1

@@ 6.2 Circuit-SAT and Weighted Circuit-SAT
Statement 3 of Theorem 6.1 implies statement 1.
@@
Statement 1 of Theorem 6.1 implies statement 4.
@@ 7.1 SETH-equivalence for Coloring
We establish that if the SETH is false, then we do get such a $q$-Coloring algorithm for parameter modulator to constant tree-depth. We are given an $n$-vertex graph $G$ and a modulator $M$ of size $m$ so that $G-M$ has tree-depth $c$. We will use Theorem 3.1 and reduce this instance to an instance of $k$-SAT with a modulator to constant tree-depth. According to Theorem 3.1, if the SETH is false, we can solve such an instance is time $2^{(1-\varepsilon)m^{\prime}}|\phi|^{O(1)}$, where $m^{\prime}$ is the size of the new modulator, for some $\varepsilon$. In particular, we will make sure that $m^{\prime}\leq(1+\frac{\varepsilon}{2})m\log q+O(1)$, and that $\phi$ has size polynomial in $n$. Then, the running time of the whole procedure will be at most $2^{(1-\varepsilon)(1+\frac{\varepsilon}{2})m\log q}n^{O(1)}=q^{(1-\varepsilon^{\prime})m}n^{O(1)}$ for some appropriate $\varepsilon^{\prime}$.
To begin our construction, we consider every vertex of $G-M$. For each such $v$, we define $q$ variables $x_{v}^{1},\ldots,x_{v}^{q}$, indicating informally that $v$ takes color $i\in[q]$. We add the clause $(x_{v}^{1}\lor\ldots\lor x_{v}^{q})$. For each $uv\in E$ with $u,v\not\in M$ we add for each $i\in[q]$ the clause ($(\neg x_{v}^{i}\lor\neg x_{u}^{i})$. It is not hard to see that the graph we have constructed so far has tree-depth at most $cq$. All the other variables we will construct will belong in the new modulator.
@@ 7.2 MaxSAT-equivalence for Max Cut
One direction is easy: given a Max Cut instance $G=(V,E)$, we can construct a CNF formula by constructing a variable $x_{v}$ for each vertex $v$ and for each edge $uv$ the clauses $(x_{u}\lor x_{v})$ and ($(\neg x_{u}\lor\neg x_{v})$. The primal graph of this formula is simply $G$, so the $(\sigma,\delta)$-hub is preserved. It is now not hard to see that the maximum number of clauses that it is possible to satisfy in the new instance is exactly $|E|+\textrm{max-cut}(G)$, so an algorithm as described in the second statement of Theorem 3.5 gives the desired algorithm for Max Cut.
@@
For each variable of $\phi$, $x_{1},\ldots,x_{n}$ we construct a vertex with the same name. We also construct a vertex $x_{0}$. For each clause $C_{j}$ of $\phi$ we construct a cycle of length $14k+$ 1, one of whose vertices is $x_{0}$, while the others are new vertices. Each edge of this cycle has weight $8k$. Number the vertices of the cycle $p_{j,1},\ldots,p_{j,4k}$, starting from a neighbor of $x_{0}$. Suppose that $C_{j}$ contains literals $\ell_{1},\ell_{2},\ldots,\ell_{k^{\prime}}$ for $k^{\prime}\leq k$. For each $i\in[k^{\prime}]$ if $\ell_{i}$ is a positive appearance of a variable $x_{r}$, then we connect $p_{j,4i+2},p_{j,4i+3}$ to $x_{r}$, otherwise if it is a negative appearance of $x_{r}$, then we connect $p_{j,4i+1},p_{j,4i+2}$ to $x_{r}$. These edges have weight $1$. Furthermore, we attach $k-k^{\prime}$ leaves to $p_{j,1}$ with edges of weight $1$. We now claim that there is an assignment satisfying $t$ clauses if and only if there is a cut of weight at least $32k^{2}m+km+t$.
@@
For the forward direction, if there is such an assignment, place all the vertices representing False variables on the same side as $x_{0}$ in the cut and all the vertices representing True variables on the other side. For each unsatisfied clause, pick a bipartition of the corresponding odd cycle so that $4k$ of the $14k+$ 1 edges of the cycle are cut and the uncut edge is incident on $x_{0}$. This contributes $32k^{2}$ from the edges of the cycle. We also obtain at least $k^{\prime}$ from the edges connecting the cycle to the variable vertices and $k-k^{\prime}$ from the attached leaves by placing each leaf to the opposite side of its neighbor, so we get $32k^{2}+k$ for each unsatisfied clause. For each satisfied clause $C_{j}$, let $\ell_{i}$ be the first literal that is set to True by the assignment. For $i^{\prime}<i$ we set $p_{j,4i^{\prime}+2},p_{j,4i^{\prime}+4}$ on the same side as $x_{0}$ and $p_{j,4i^{\prime}+1},p_{j,4i^{\prime}+3}$ on the other side; while for $i^{\prime}>i$ we use the opposite partition. We now have that the vertex preceding $p_{j,4i+1}$ and the vertex after $p_{j,4i+4}$ in the cycle are both on the side of $x_{0}$. For the four remaining vertices, if $\ell_{i}$ is a variable that appears positive, we set $p_{j,4i+2},p_{j,4i+3}$ on the same side as $x_{0}$, cutting both edges connecting these vertices to the corresponding variable vertex, and $p_{j,4i+1},p_{j,4i+4}$ on the other side; while if $\ell_{i}$ is a variable that appears negative, we set $p_{j,4i+1},p_{j,4i+2},p_{j,4i+4}$ on the side opposite $x_{0}$ and $p_{j,4i+3}$ on the side of $x_{0}$, again cutting both edges to the corresponding variable vertex. We now observe that for a satisfied clause we have cut $132k^{2}+k+$ 1 edges, obtaining the bound on the total size of the cut.
@@
For the converse direction we observe that an optimal cut must cut $4k$ of the $14k+$ 1 edges of the odd cycle constructed for each clause. Indeed, it is not possible to cut more (since the cycle is odd), while if we leave two edges uncut, we may replace the solution by a solution that cuts $4k$ of the edges. This will add $8k$ to the total weight, more than making up for the at most $2k$ lost edges connecting the cycle to the rest of the graph. Thus, for each cycle we obtain $32k^{2}$ from the edges inside the cycle. Furthermore, we can obtain at most $k+1$ from the remaining edges incident on the cycle. To see this, note that we have added $k^{\prime}$ pairs of edges (one pair for each literal of the clause) which are adjacent to consecutive vertices in the cycle. If we cut both edges in a pair, an edge of the cycle is uncut, so there is at most one pair of cut edges and from each other pair we cut at most one edge. Taking into account the $k-k^{\prime}$ incident on leaves (which are always cut), we have that each cycle contributes at most $132k^{2}+k+$ 1 to the cut. Hence, there must be at least $t$ cycles contributing exactly this much, which we claim correspond to satisfied clauses by the assignment we extract by setting to False variables whose corresponding vertices are on the same side as $x_{0}$. Indeed, take a cycle representing $C_{j}$ and suppose that $C_{j}$ is not satisfied by this assignment. However, there exists a literal $\ell_{i}$ in $C_{j}$ such that we cut both edges added for $\ell_{i}$. If $\ell_{i}$ is positive, then $p_{j,4i+2},p_{j,4i+3}$ have a common neighbor. If $C_{j}$ is not satisfied, this neighbor must be on the same side as $x_{0}$, so $p_{j,4i+2},p_{j,4i+3}$ are on the other side. It is now not hard to see via a parity argument that two edges of the cycle must be uncut, contradiction. The reasoning is similar if $\ell_{i}$ is negative, as then $p_{j,4i+1},p_{j,4i+2}$ must have a common neighbor, which must be on the side not containing $x_{0}$ (otherwise $C_{j}$ is satisfied), so $p_{j,4i+1},p_{j,4i+2}$ are on the same side as $x_{0}$, and again two edges are uncut. Hence, a cut of the specified size leads to an assignment satisfying at least $t$ clauses.
@@ 7.3 Coloring and Pathwidth Modulators
For all $q\geq 3$ we have the following. There is an $\varepsilon>0$ such that for all $c>0$ there is an algorithm that takes as input an $n$-vertex graph $G$ and a modulator $M$ of size $m$ such that $G-M$ has pathwidth at most $c$ and decides if $G$ is $q$-colorable in time $(q-\varepsilon)^{m}n^{O(1)}$ if and only if the LD-C-SETH is false. The same is true if we replace “pathwidth $c$” with “tree-depth $c\log n$” in the preceding statement.
@@
The first claim readily follows from the proof of Theorem 7.1. Recall that in that proof we showed how to transform an instance $G$ of $q$-Coloring with a modulator $M$ to tree-depth $c$ to an instance of SAT with a modulator of size essentially $|M|\log q$, while ensuring that in the new instance the removal of the modulator leaves a graph of tree-depth at most $cq$. We run the same reduction, except now the tree-depth is at most $cq\log n$. One of the statements of Theorem 4.1 states that we can avoid brute-force for SAT parameterized by a modulator to logarithmic tree-depth if and only if the LD-C-SETH is false.
@@
We are given a CNF formula $\phi$ and a modulator $M$ of size $m$ such that $\phi-M$ has pathwidth $c$. Using Section A.2 we can assume that $\phi$ is a 3-CNF formula. Select integers $\gamma,\rho$ such that we have the following:
@@
We first reduce the arity of the CSP instance to $2$. For each constraint $C$ involving more than $2$ and at most $3\gamma$ variables, let $\mathcal{S}_{C}$ be the set of acceptable assignments, which has cardinality at most $q^{3\gamma}=O(1)$. We construct $12|\mathcal{S}_{C}|+$ 1 new variables and connect them in an odd cycle with constraints of arity $2$, ensuring that all the variables which are involved in a constraint take distinct values, that $|\mathcal{S}_{C}|$ variables, which are pairwise non-adjacent, have possible values $\{1,2,3\}$, and that all the others have possible values $\{1,2\}$. The intuition here is that we have formed an odd cycle and we can use colors $\{1,2,3\}$ to color it. Since the cycle is odd, we must use color $3$ somewhere, and if we use color $3$ once we can 2-color the rest. The idea then is to associate each of the $|\mathcal{S}_{C}|$ variables that have three possible values with a satisfying assignment to $C$ and add constraints that ensure that if this variable takes color $3$ then the variables of $C$ must agree with the corresponding assignment. For this, for each of the $|\mathcal{S}_{C}|$ variables of the cycle that represents some assignment $\sigma$ we add constraints of arity $2$ involving this variable and each variable of $C$. The constraints state that if the variable of the cycle has value $1$ or $2$, then everything is acceptable, otherwise the other variable must agree with $\sigma$. We do this exhaustively and it is not hard to see that the answer is preserved. We have to argue that the pathwidth of the primal graph (after removing the modulator) is almost preserved. Indeed, for each constraint $C$, find a bag that contains all its variables (outside of $M$), and after this bag add a sequence of copies of it, where we add a path decomposition of the cycle we have constructed. Repeating this will not add to the pathwidth more than a small additive constant.
@@
We will show reductions in both directions from Ann-s-t-Non-Reach which was established to be equivalent to 2SatB-SETH in Section 5. As a reminder, in Ann-s-t-Non-Reach we are given a DAG with its arcs annotated from a set of $m$ boolean variables. We are asked to select an assignment to the variables so that in the DAG resulting from keeping only non-annotated arcs or arcs whose annotation is set to True by the assignment there is no $s\to t$ path, for two given vertices $s,t$. We have shown that 2SatB-SETH is equivalent to assuming that it is impossible to solve Ann-s-t-Non-Reach without considering essentially all assignments to the $m$ boolean variables.
@@
Suppose that there is an algorithm for $k$-Neighborhood-Cut running in time $N^{k-\varepsilon}$ for some fixed $k\geq 2,\varepsilon>0$. We can now solve Ann-s-t-Non-Reach by running the reduction described above and then solving $k$-Neighborhood-Cut. The running time will be dominated by the second part, which would take $N^{k-\varepsilon}=2^{\frac{m}{k}(k-\varepsilon)}n^{O(1)}=2^{(1-\frac{\varepsilon}{k})m}n^{O(1)}$, which would falsify the 2SatB-SETH by the results of Section 5.
@@ 7.5 Degenerate Deletion
For the direction that Circuit-SETH implies hardness for our problem we essentially reuse the reduction of [36], which relies on a previous reduction of [42], and observe that these reductions from the standard W[P]-complete problem we dealt with in Theorem 6.1 actually preserve the value of $k$ exactly and only increase the size of the instance by a constant factor, therefore the lower bound is preserved as well. For the other direction we will perform a direct reduction to Circuit-SAT.
Let us give the relevant details. For the first direction, we start with an instance of monotone weight-$k$ circuit satisfiability. By Theorem 6.1, if we can solve such an instance in time $s^{(1-\varepsilon)k}$ for monotone circuits of $n$ inputs and size $s$, assuming $k$ is a large enough constant, we will falsify the Circuit-SETH. As shown by Szeider [42], there is a simple reduction from this problem to the Cyclic Monotone Circuit Activation problem. In the latter problem we are given a monotone boolean circuit which is a general digraph (not a DAG) and are asked to select $k$ initial gates to activate (set to True). From that point on at each round a gate becomes True if at least one of its in-neighbors is True for disjunction ($\lor$) gates and if all its in-neighbors are True, for conjunction ($\land$) gates. The question is whether there is a selection of $k$ initial gates that eventually makes everything True.

### v2

@@ 6.2 Circuit-SAT and Weighted Circuit-SAT
Statement 3 of Theorem˜6.1 implies statement 1.
@@
Statement 1 of Theorem˜6.1 implies statement 4.
@@ 7.1 SETH-equivalence for Coloring
We establish that if the SETH is false, then we do get such a $q$-Coloring algorithm for parameter modulator to constant tree-depth. We are given an $n$-vertex graph $G$ and a modulator $M$ of size $m$ so that GMG-M has tree-depth $c$. We will use Theorem˜3.1 and reduce this instance to an instance of $k$-SAT with a modulator to constant tree-depth. According to Theorem˜3.1, if the SETH is false, we can solve such an instance is time $2^{(1-\varepsilon)m^{\prime}}|\phi|^{O(1)}$, where $m^{\prime}$ is the size of the new modulator, for some $\varepsilon$. In particular, we will make sure that $m^{\prime}\leq(1+\frac{\varepsilon}{2})m\log q+O(1)$, and that $\phi$ has size polynomial in $n$. Then, the running time of the whole procedure will be at most $2^{(1-\varepsilon)(1+\frac{\varepsilon}{2})m\log q}n^{O(1)}=q^{(1-\varepsilon^{\prime})m}n^{O(1)}$ for some appropriate $\varepsilon^{\prime}$.
To begin our construction, we consider every vertex of GMG-M. For each such $v$, we define $q$ variables $x_{v}^{1},\ldots,x_{v}^{q}$, indicating informally that $v$ takes color $i\in[q]$. We add the clause $(x_{v}^{1}\lor\ldots\lor x_{v}^{q})$. For each $uv\in E$ with $u,v\not\in M$ we add for each $i\in[q]$ the clause $(\neg x_{v}^{i}\lor\neg x_{u}^{i})$. It is not hard to see that the graph we have constructed so far has tree-depth at most $cq$. All the other variables we will construct will belong in the new modulator.
@@ 7.2 MaxSAT-equivalence for Max Cut
One direction is easy: given a Max Cut instance $G=(V,E)$, we can construct a CNF formula by constructing a variable $x_{v}$ for each vertex $v$ and for each edge $uv$ the clauses $(x_{u}\lor x_{v})$ and $(\neg x_{u}\lor\neg x_{v})$. The primal graph of this formula is simply $G$, so the $(\sigma,\delta)$-hub is preserved. It is now not hard to see that the maximum number of clauses that it is possible to satisfy in the new instance is exactly $|E|+\textrm{max-cut}(G)$, so an algorithm as described in the second statement of Theorem˜3.5 gives the desired algorithm for Max Cut.
@@
For each variable of $\phi$, $x_{1},\ldots,x_{n}$ we construct a vertex with the same name. We also construct a vertex $x_{0}$. For each clause $C_{j}$ of $\phi$ we construct a cycle of length $4k+1$, one of whose vertices is $x_{0}$, while the others are new vertices. Each edge of this cycle has weight $8k$. Number the vertices of the cycle $p_{j,1},\ldots,p_{j,4k}$, starting from a neighbor of $x_{0}$. Suppose that $C_{j}$ contains literals $\ell_{1},\ell_{2},\ldots,\ell_{k^{\prime}}$ for $k^{\prime}\leq k$. For each $i\in[k^{\prime}]$ if $\ell_{i}$ is a positive appearance of a variable $x_{r}$, then we connect $p_{j,4i+2},p_{j,4i+3}$ to $x_{r}$, otherwise if it is a negative appearance of $x_{r}$, then we connect $p_{j,4i+1},p_{j,4i+2}$ to $x_{r}$. These edges have weight $1$. Furthermore, we attach $k-k^{\prime}$ leaves to $p_{j,1}$ with edges of weight $1$. We now claim that there is an assignment satisfying $t$ clauses if and only if there is a cut of weight at least $32k^{2}m+km+t$.
@@
For the forward direction, if there is such an assignment, place all the vertices representing False variables on the same side as $x_{0}$ in the cut and all the vertices representing True variables on the other side. For each unsatisfied clause, pick a bipartition of the corresponding odd cycle so that $4k$ of the $4k+1$ edges of the cycle are cut and the uncut edge is incident on $x_{0}$. This contributes $32k^{2}$ from the edges of the cycle. We also obtain at least $k^{\prime}$ from the edges connecting the cycle to the variable vertices and $k-k^{\prime}$ from the attached leaves by placing each leaf to the opposite side of its neighbor, so we get $32k^{2}+k$ for each unsatisfied clause. For each satisfied clause $C_{j}$, let $\ell_{i}$ be the first literal that is set to True by the assignment. For $i^{\prime}<i$ we set $p_{j,4i^{\prime}+2},p_{j,4i^{\prime}+4}$ on the same side as $x_{0}$ and $p_{j,4i^{\prime}+1},p_{j,4i^{\prime}+3}$ on the other side; while for $i^{\prime}>i$ we use the opposite partition. We now have that the vertex preceding $p_{j,4i+1}$ and the vertex after $p_{j,4i+4}$ in the cycle are both on the side of $x_{0}$. For the four remaining vertices, if $\ell_{i}$ is a variable that appears positive, we set $p_{j,4i+2},p_{j,4i+3}$ on the same side as $x_{0}$, cutting both edges connecting these vertices to the corresponding variable vertex, and $p_{j,4i+1},p_{j,4i+4}$ on the other side; while if $\ell_{i}$ is a variable that appears negative, we set $p_{j,4i+1},p_{j,4i+2},p_{j,4i+4}$ on the side opposite $x_{0}$ and $p_{j,4i+3}$ on the side of $x_{0}$, again cutting both edges to the corresponding variable vertex. We now observe that for a satisfied clause we have cut $32k^{2}+k+1$ edges, obtaining the bound on the total size of the cut.
@@
For the converse direction we observe that an optimal cut must cut $4k$ of the $4k+1$ edges of the odd cycle constructed for each clause. Indeed, it is not possible to cut more (since the cycle is odd), while if we leave two edges uncut, we may replace the solution by a solution that cuts $4k$ of the edges. This will add $8k$ to the total weight, more than making up for the at most $2k$ lost edges connecting the cycle to the rest of the graph. Thus, for each cycle we obtain $32k^{2}$ from the edges inside the cycle. Furthermore, we can obtain at most $k+1$ from the remaining edges incident on the cycle. To see this, note that we have added $k^{\prime}$ pairs of edges (one pair for each literal of the clause) which are adjacent to consecutive vertices in the cycle. If we cut both edges in a pair, an edge of the cycle is uncut, so there is at most one pair of cut edges and from each other pair we cut at most one edge. Taking into account the $k-k^{\prime}$ incident on leaves (which are always cut), we have that each cycle contributes at most $32k^{2}+k+1$ to the cut. Hence, there must be at least $t$ cycles contributing exactly this much, which we claim correspond to satisfied clauses by the assignment we extract by setting to False variables whose corresponding vertices are on the same side as $x_{0}$. Indeed, take a cycle representing $C_{j}$ and suppose that $C_{j}$ is not satisfied by this assignment. However, there exists a literal $\ell_{i}$ in $C_{j}$ such that we cut both edges added for $\ell_{i}$. If $\ell_{i}$ is positive, then $p_{j,4i+2},p_{j,4i+3}$ have a common neighbor. If $C_{j}$ is not satisfied, this neighbor must be on the same side as $x_{0}$, so $p_{j,4i+2},p_{j,4i+3}$ are on the other side. It is now not hard to see via a parity argument that two edges of the cycle must be uncut, contradiction. The reasoning is similar if $\ell_{i}$ is negative, as then $p_{j,4i+1},p_{j,4i+2}$ must have a common neighbor, which must be on the side not containing $x_{0}$ (otherwise $C_{j}$ is satisfied), so $p_{j,4i+1},p_{j,4i+2}$ are on the same side as $x_{0}$, and again two edges are uncut. Hence, a cut of the specified size leads to an assignment satisfying at least $t$ clauses.
@@ 7.3 Coloring and Pathwidth Modulators
For all $q\geq 3$ we have the following. There is an $\varepsilon>0$ such that for all $c>0$ there is an algorithm that takes as input an $n$-vertex graph $G$ and a modulator $M$ of size $m$ such that GMG-M has pathwidth at most $c$ and decides if $G$ is $q$-colorable in time $(q-\varepsilon)^{m}n^{O(1)}$ if and only if the LD-C-SETH is false. The same is true if we replace “pathwidth cc” with “tree-depth $c\log n$” in the preceding statement.
@@
The first claim readily follows from the proof of Theorem˜7.1. Recall that in that proof we showed how to transform an instance $G$ of $q$-Coloring with a modulator $M$ to tree-depth $c$ to an instance of SAT with a modulator of size essentially $|M|\log q$, while ensuring that in the new instance the removal of the modulator leaves a graph of tree-depth at most $cq$. We run the same reduction, except now the tree-depth is at most $cq\log n$. One of the statements of Theorem˜4.1 states that we can avoid brute-force for SAT parameterized by a modulator to logarithmic tree-depth if and only if the LD-C-SETH is false.
@@
We are given a CNF formula $\phi$ and a modulator $M$ of size $m$ such that $\phi-M$ has pathwidth $c$. Using Section˜A.2 we can assume that $\phi$ is a 3-CNF formula. Select integers $\gamma,\rho$ such that we have the following:
@@
We first reduce the arity of the CSP instance to $2$. For each constraint $C$ involving more than $2$ and at most $3\gamma$ variables, let $\mathcal{S}_{C}$ be the set of acceptable assignments, which has cardinality at most $q^{3\gamma}=O(1)$. We construct $2|\mathcal{S}_{C}|+1$ new variables and connect them in an odd cycle with constraints of arity $2$, ensuring that all the variables which are involved in a constraint take distinct values, that $|\mathcal{S}_{C}|$ variables, which are pairwise non-adjacent, have possible values $\{1,2,3\}$, and that all the others have possible values $\{1,2\}$. The intuition here is that we have formed an odd cycle and we can use colors $\{1,2,3\}$ to color it. Since the cycle is odd, we must use color $3$ somewhere, and if we use color $3$ once we can 2-color the rest. The idea then is to associate each of the $|\mathcal{S}_{C}|$ variables that have three possible values with a satisfying assignment to $C$ and add constraints that ensure that if this variable takes color $3$ then the variables of $C$ must agree with the corresponding assignment. For this, for each of the $|\mathcal{S}_{C}|$ variables of the cycle that represents some assignment $\sigma$ we add constraints of arity $2$ involving this variable and each variable of $C$. The constraints state that if the variable of the cycle has value $1$ or $2$, then everything is acceptable, otherwise the other variable must agree with $\sigma$. We do this exhaustively and it is not hard to see that the answer is preserved. We have to argue that the pathwidth of the primal graph (after removing the modulator) is almost preserved. Indeed, for each constraint $C$, find a bag that contains all its variables (outside of $M$), and after this bag add a sequence of copies of it, where we add a path decomposition of the cycle we have constructed. Repeating this will not add to the pathwidth more than a small additive constant.
@@
We will show reductions in both directions from Ann-s-t-Non-Reach which was established to be equivalent to 2SatB-SETH in Section˜5. As a reminder, in Ann-s-t-Non-Reach we are given a DAG with its arcs annotated from a set of $m$ boolean variables. We are asked to select an assignment to the variables so that in the DAG resulting from keeping only non-annotated arcs or arcs whose annotation is set to True by the assignment there is no $s\to t$ path, for two given vertices $s,t$. We have shown that 2SatB-SETH is equivalent to assuming that it is impossible to solve Ann-s-t-Non-Reach without considering essentially all assignments to the $m$ boolean variables.
@@
Suppose that there is an algorithm for $k$-Neighborhood-Cut running in time $N^{k-\varepsilon}$ for some fixed $k\geq 2,\varepsilon>0$. We can now solve Ann-s-t-Non-Reach by running the reduction described above and then solving $k$-Neighborhood-Cut. The running time will be dominated by the second part, which would take $N^{k-\varepsilon}=2^{\frac{m}{k}(k-\varepsilon)}n^{O(1)}=2^{(1-\frac{\varepsilon}{k})m}n^{O(1)}$, which would falsify the 2SatB-SETH by the results of Section˜5.
@@ 7.5 Degenerate Deletion
For the direction that Circuit-SETH implies hardness for our problem we essentially reuse the reduction of [41], which relies on a previous reduction of [48], and observe that these reductions from the standard W[P]-complete problem we dealt with in Theorem˜6.1 actually preserve the value of $k$ exactly and only increase the size of the instance by a constant factor, therefore the lower bound is preserved as well. For the other direction we will perform a direct reduction to Circuit-SAT.
Let us give the relevant details. For the first direction, we start with an instance of monotone weight-$k$ circuit satisfiability. By Theorem˜6.1, if we can solve such an instance in time $s^{(1-\varepsilon)k}$ for monotone circuits of $n$ inputs and size $s$, assuming $k$ is a large enough constant, we will falsify the Circuit-SETH. As shown by Szeider [48], there is a simple reduction from this problem to the Cyclic Monotone Circuit Activation problem. In the latter problem we are given a monotone boolean circuit which is a general digraph (not a DAG) and are asked to select $k$ initial gates to activate (set to True). From that point on at each round a gate becomes True if at least one of its in-neighbors is True for disjunction ($\lor$) gates and if all its in-neighbors are True, for conjunction ($\land$) gates. The question is whether there is a selection of $k$ initial gates that eventually makes everything True.